# Regular triangle unions with maximal number of sides

### v3

@@
How to prove the stated property? Since we go along the common boundary counter clock-wise, each leg has a direction. Drawing a full line through its endpoints, we can talk about right and left side of it. The same applies to any chosen edge. All the remaining edges split into lying to the right, to the left of it, or collinear. Also, for any two points on the same leg, we can say which one is behind and which one is ahead the other. Now, fix any number in the produced seq. of right edges. All the successive edges are either collinear (those with the same number), or lie to the right of it. Suppose, a subseq. abaaba belongs to such a seq. Let legs $a$ and $b$ intersect at a point $P$. In the process of recording right edges, we pass from leg $a$ to leg $b$, and then to leg $a$ again. Stop here. The point $P$ now lies behind. It is no longer possible to get beck to leg $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
Hence, $\mathcal{L}$ and all such seq. are exactly what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq. by definition sharir . The maximal length of $(2,n)$-DS is known to be 2n12n-1.
@@
The proof of this fact is done by induction. Suppose, it is true for n1n-1. Consider a $(2,n)$-DS seq. $\mathcal{L}$. If not all symbols $1$ through $n$ appear, it is then a $(2,m)$-DS seq. for some $m<n$. Otherwise, reading left-to-right, assume that symbols appear in their natural order: $1,2,\ldots,n$ (one can always rename if needed). We claim that symbol $n$ appears in $\mathcal{L}$ only once. Assuming the opposite, let $\mathcal{L}$ contain a subseq. npnnpn for some $p\leq n-1$. Due to natural order convention, $\mathcal{L}$ must also contain a subseq. pnpnpnpn – a contradiction. Removing $n$ and concatenating two symbols which surround it in case they are equal, we get a (2,n1)(2,n-1)-DS seq. of a maximal length 2n32n-3. Thus, $\mathcal{L}$ has a length at most 2n3+22n-3+2. Note also that in case $(2,n)$-DS seq. is maximal, the last symbol to appear must always be surrounded by two identical symbols.
@@
Let us return to our edge count. Take the same underlined fragment, but start from the other end. That is, write down numbers of $\triangle$ whose LEFT edge contribute to the boundary. This way, seq. $18781654613231$ is obtained. It is again a $(2,n)$-DS seq. Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue edges were counted twice. We perform the same calculation for all three blue vertex pairs. Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of $M$-gon, then
@@
The inequality $e(n)<12n$ also implies the following. Take a regular $\bigcup$ of $n$ $\triangle$ with $e(n)$ edges. Choose $\triangle$ which contributes least to the common boundary. Its toll is thus $\leq 11$ edges. Remove this $\triangle$ from the $\bigcup$. Each edge not adjacent to its vertex reduces an edge count by $1$, while two edges at any vertex reduce this by $3$. Therefore, $e(n-1)\geq e(n)-14$.
The idea to separate left and right edges between two same-colored vertices was inspired by the proof of the aforementioned Zone Theorem given by Edelsbrunner at al. in edelsbrunner . In the latter paper though, ‘right’ and ‘left’ correspond to two directions along the same ray. Among several existing proofs of the Zone Theorem, it is the one which essentially employs $(2,n)$-DS sequences.
What this result claims? Consider $\mathcal{L}$, a simple arrangement of $n$ lines in the affine plane. The set $\mathbb{R}^{2}\setminus\mathcal{L}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges. The size of a face is the number of its edges. Given a line $\ell$. If $\mathcal{L}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses. The complexity of the zone of $\ell$ is the sum of sizes of all faces in the zone. This notion was first introduced by Chazelle at al. in zone , where an upper bound $6n$ for complexity was given. DS seq. approach yields 6n46n-4 (see edelsbrunner , p. 334, also sharir , p. 125). In horizon , Bern at al. strengthened this to $\lfloor 5.5n\rfloor-1$. Rom Pinchasi pinchasi later improved it to $\lfloor 5.5n\rfloor-3$ pinchasi . A construction in Fig. 13 (taken from horizon , just slightly enhanced) shows that the latter bound is almost optimal. Note the similarity with construction in Fig. 11.
Figure 13: An arrangement of $11$ lines; a complexity of the zone of blue line (colored in green) is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
@@
The final insight comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq. must be non-maximal. In terms of the upper bound, this only settles the case $n=7$, and reduces it by $1$ for $15\leq n\leq 18$. This seems a feeble contribution. However, such deeper analysis also gives an algorithm to construct a seq. of length $\mathscr{U}(n)$ which does not violate combinatorics of pseudolines.
For the starters, let us return to the example $\mathcal{L}_{1}=123214546417871$. Its length is $2\cdot 8-1$. So, it is maximal ($\mathbf{M}$ for short). Now, shift our attention to the cyan triangle. Or, better, the overlined fragment in (1). The produced seq. is also M. We argue that this extension to $\mathcal{L}_{1}$ can be reproduced from $\mathcal{L}_{1}$ itself. Indeed, starting from $\mathcal{L}_{1}$, remove all (but the last) $1$’s and place $2$ in the end. This results in $232454647871\ldots 2$. For this to be $\mathbf{M}$, dots should be replaced by two numbers: $\mathcal{L}_{2}=232454647871ab2$. We claim that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$. The argument behind this will now certainly look familiar. Let us methodically remove the last (reading left to right) symbol to appear, and merge its two neighbours. Such merge must always occur:
@@
This gives the only possible reconstruction: $\mathcal{L}_{2}=232454647871742$. The overlined fragment in (1) confirms it. In general, a completely identical argument proves the following statement.
Lemma.
Given a prefix $\mathcal{P}=1\ldots 2\ldots 3\ldots n$ and the terminal number $1$. If $\mathcal{P}1$ is a $(2,n)$-DS seq., there exists the unique $\mathcal{Q}$ such that $\mathcal{P}\mathcal{Q}1=\mathscr{S}(Q)$ is maximal.
This allows to iterate the procedure (reconstructed symbols are printed in red):
@@
This exactly (!) corresponds to actual seq. extracted from (1). So far, all of them are $\mathbf{M}$. We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}_{1}$ by increasing all terms by 1. As we will soon see, this is a general rule.
Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences extracted from (1). For a while, $\mathbf{M}$ seq. continue. However, at a step $17$ (double underline), a nonmaximal sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs. Then it is followed with a $\mathbf{M}$ $2345436768631{\color[rgb]{1,0,0}3}2$. Next $2$ seq. are also $\mathbf{M}$. Then comes another shorther $56768612321645$ (length $14$), after that comes $\mathbf{M}$ $6768612321645{\color[rgb]{1,0,0}4}6$. Finally, after two more steps, everything closes up at the initial $1232145464178{\color[rgb]{1,0,0}7}1$. The conclusion: process almost always runs smoothly (and the next seq. is deterministic), only during $24$ steps two glitches occur. The total number of edges in the $\bigcup$ equals the number of red symbols in this sequence of sequences (each of length 15 or 14), plus $6\cdot 8$.
If the reader has already decided that this is too confusing to follow, this is true. Seeing such an arbitrary-looking collection of numbers in (3), it is impossible to fathom what is actually happening. In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of regular $(n+1)$-gons (first discovered by D. Roselle roselle ; see also sharir , p.250-232), we will see that… almost nothing is happening!
To witness this, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$. First, divide it into $9$ decreasing portions as follows: | 1| 2| 3| 43| 53| 6| 76| 86321| 91||\,1|\,2\,|\,3\,|\,43\,|\,53\,|\,6\,|\,76\,|\,86321\,|\,91\,|. Given regular $(n+1)$-gon. For each portion starting from $i$, connect vertices with these particular labels to the vertex $i+1$. This way Fig. 14 is obtained.
@@
It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after redefining or reordering initial set of symbols).
In our case, this triangulation convention is used for the natural order: $1,2,\ldots,n$. If the order is $i,i+1,\ldots i-1$, we adopt the following rule. First, rotate symbols cyclically, so that the order is $1,2,\ldots,n$. So, for $\mathscr{S}(\mathcal{L}_{1})=232454647871742$ one gets $121343536768631$. Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to position $i$.
Let as apply this rule for visualising actual seq. extracted from (1). The result is presented as Fig. 15. It is surprising and relieving at the same time: nothing changes going from seq. 1 to seq. 2, except that the distinguished vertex moved one position. In light of such triangulation correspondence, Lemma become a tautology. Given a triangulation of a regular $n$-gon. Then it tells that if one removes all edges ending at one particular vertex, then there is only one way to complete the triangulation by adding edges starting from this particular vertex.
Such pattern continues up to seq. 8. Going to seq. 9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction. This becomes clear if given some thought.
Now, what happens when a glitch (a shorter seq.) occurs? How it is succeeded by a $\mathbf{M}$ seq.? The same Fig. 15 (in the traditions of Intelligencer) might be considered as a proof without words of the following fact: at step $16\rightarrow 17$ (distinguished is vertex 8), one edge ending with $8$ is removed (at the same time, since the vertex changes from $8$ to $1$, a one-step rotation in positive direction occurs). At step $17\rightarrow 18$ (vertex 1) and edge with endpoint $1$ is added. After two more moves, another such incident takes place: one edge is removed, a new edge is reinstated during the next move.
We come to understand that the whole process runs as follows. Start from a distinguished vertex in a triangulation of a regular $(n+1)$-gon. Perform $3n$ moves. An $i$th move amounts to removing diagonals at this vertex; adding other diagonals (at the same vertex); counting their number $t_{i}$; switching to the next neighbour. After the whole operation, vertex n1n-1 is reached. The goal is to obtain the same triangulation, only rotated $3$ turns in the negative direction. The total combinatorial length of such $\triangle$ $\bigcup$ is then $\sum_{i=1}^{3n}t_{i}+6n$.
@@
Figure 15: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 12; 1-9 also correspond exactly to seq. in (3)
If no deletions are necessary, it means that the triangulation coincides with itself if rotated by 3-3 moves. So, $\alpha(n)$ equals $6$ times the total number of edges in the triangulation, that is, 6(n+1+n2)=12n66(n+1+n-2)=12n-6, minus the sum of number of edges at three consecutive vertices. For $n+1=6$ and $9$ we see that this sum is equal to $9$ and $10$, respectively, independently of the start vertex. As expected, this gives $12\cdot 5-6-9=45$ and $12\cdot 8-6-10=80$ edges of the $\bigcup$.
@@
This section is very sketchy, mostly referring to the monograph of Sharir and Agarwal sharir . Take $7$ linear univariate functions defined on the real line. Label them $1$ through $7$. Let $f$ be defined as the minimum of these. If, while taking the minimal value, we jump from one line to the other, it is obviously no longer possible to go back to the abandoned one. Therefore, the support sequence (telling which function provides the minimum as one progresses along the $x$-axis) for $f$ consist of at most $7$ symbols.
However, if each linear function is defined on its own interval so that their $\bigcup$ is also an interval, the support sequence can be longer. This is clear. What is astonishing that it can be much, much, longer! As an example, consider $7$ segments as in Fig. 16.
@@
The support seq. $1213145416164653737572767$ has length $25$. What properties does such seq. have? Similarly as with $\mathcal{L}$, no two adjacent symbols are equal. At the same time, we immediately notice its subseq. like $1313$, $1414$, $3535$, etc. This is a major distinction. We leave it as an exercise for th reader to prove that though subseq. abababab may be present, subseq. of the for ababaababa are not to be found. This is exactly the definition of a $(3,n)$-Davenport-Schintzel seq.
@@
The foundational part of sharir (p. 12-42) is devoted to investigating $(3,n)$-DS seq. The story how their true complexity unravelled is inspiring. The final answer – the maximal length of such seq. is $2n\alpha(n)+O(n)$, a result by Seth Pettie pettie – tells little about the story. In the preface to the book is is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
@@
Here is another major achievement. Consider a $(3,n)$-DS, now of maximal possible length $27$ (see seq. A002004 in OEIS): $121314151616543272737475767$. It can be realized as a lower envelope of $7$ pseudosegments (connected parts of $7$ pseudolines). Can it be realized as a lower envelope of $7$ segments? In other words, can pseudosegment construction be stretched? I do not know. What is certain that there is little hope (and most likely, untrue; see sharir , p. 112) that a certain maximal $(3,n)$-DS seq. for large $n$ can serve as an envelope of $n$ segments. However, Wiernik and Sharir proved that there exist $(3,n)$-DS seq. whose length asymptotically differs by a factor of at most $2$ from the length of longest such seq. and which does serve as lower envelope of $n$ segments. Speaking philosophically, this shows that the Ackermann function, first discovered in logic and computability analysis, is intricately present in nature, if we assume that straight segments are. Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is something what carries mathematics well outside the domain of natural sciences.
@@
To finish the construction, consider Fig. 16 once again. If we imagine each segment completed to a triangle by two nearly vertical legs, we clearly can get a simple polygon. Thus, finding $i(n)$ rests on other fundamental problems about unions. The following problem might serve a good training field:
Problem.
@@
During the Journées Arithmétiques conference held in Vilnius, Lithuania in 2011, Artūras Dubickas gave a talk. Andrzej Schinzel, then in the audience, noted the question posed by the speaker. Later he wrote a note containing the solution (Fig. 17). That is how an autograph of a famous mathematician landed in a country-side of Northern Lithuania. Though it is a bit of a stretch, but numbers $n=1,2,5,8$ which give rise to perfect trangle unions are also in the Fibonacci sequence.
Figure 17: A handwritten note made by Andrzej Schinzel in Vilnius, 2011

### v4

@@
How to prove the stated property? Since we go along the boundary counter clockwise, each leg has a direction. Drawing a full line through its endpoints, we can talk about right and left sides of it. The same applies to any chosen edge. All the remaining edges split into lying to the right, to the left, or collinear with it. Also, for any two points on the same leg, we can say which one is behind and which one is ahead the other. Now, fix any number in the produced seq. (like $\mathcal{L}$) of right edges. All the successive edges are either collinear (those with the same number), or lie to the right of it. Suppose, a subseq. abaaba belongs to such a seq. Let legs $a$ and $b$ intersect at a point $P$. In the process of recording right edges, we pass from leg $a$ to leg $b$, and then to leg $a$ again. Stop here. The point $P$ now lies behind. It is no longer possible to get back unto $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
@@
Hence, $\mathcal{L}$ and all seq. of right edges between two same-colored vertices of a fixed $\triangle$ are what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq. by very definition sharir . The maximal length of $(2,n)$-DS is known to be 2n12n-1. The proof is done by induction. Suppose, it is true for n1n-1. Consider a $(2,n)$-DS seq. $\mathcal{J}$. If not all symbols $1$ through $n$ appear, it is then a $(2,m)$-DS seq. for some $m<n$. Otherwise, reading left-to-right, assume that symbols appear in their natural order: $1,2,\ldots,n$ (one can always rename if needed). We claim that symbol $n$ appears in $\mathcal{J}$ only once. Assuming the opposite, let $\mathcal{J}$ contain a subseq. npnnpn for some $p\leq n-1$. Due to a natural order convention, $\mathcal{J}$ must also contain a subseq. pnpnpnpn – a contradiction. Removing $n$ and concatenating two symbols which surround it in case they are equal, we get a (2,n1)(2,n-1)-DS seq. of a maximal length 2n32n-3. Thus, $\mathcal{J}$ has a length at most 2n3+22n-3+2. Note also that in case $(2,n)$-DS seq. is maximal, the last symbol to appear must always be surrounded by two identical symbols.
@@
Let us return to our edge count. Take the same underlined fragment, but start from the other end. That is, write down numbers of $\triangle$ whose LEFT edge contribute to the boundary. This way, seq. $18781654613231$ is obtained. It is again a $(2,n)$-DS seq. (non-maximal this time). Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue ones were counted twice. We perform the same calculation for all three blue vertex pairs. Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of a $M$-gon, then
@@
The inequality $e(n)<12n$ also implies the following. Take a regular $\bigcup$ of $n$ $\triangle$ with $e(n)$ edges. Choose $\triangle$ which contributes least to the common boundary. Its toll is thus $\leq 11$ edges. Remove this $\triangle$ from the $\bigcup$. Each edge not adjacent to its vertex reduces an edge count by (at most) $1$, while two edges at any vertex reduce this number by (at most) $3$. Therefore, $e(n-1)\geq e(n)-14$.
The idea to separate left and right edges between two same-colored vertices was inspired by the proof of the aforementioned Zone Theorem given by Edelsbrunner at al. in edelsbrunner . In the latter paper though, ‘right’ and ‘left’ correspond to two directions along the same ray.
Figure 12: An arrangement of $11$ lines; a complexity of the zone (colored in green) of a blue line is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
Among several existing proofs of the Zone Theorem, it is the one which essentially employs $(2,n)$-DS sequences.
What it claims? Consider $\mathcal{K}$, a simple arrangement of $n$ lines in the affine plane. The set $\mathbb{R}^{2}\setminus\mathcal{K}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges. The size of a face is the number of its edges. Given a line $\ell$. If $\mathcal{K}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses. The complexity of the zone of $\ell$ is the sum of sizes of all faces in the zone. This notion was first introduced by Chazelle at al. in zone , where an upper bound $6n$ for the complexity was given. DS seq. approach yields 6n46n-4 (see edelsbrunner , p. 334, also sharir , p. 125). In horizon , Bern at al. strengthened this to $\lfloor 5.5n\rfloor-1$. Rom Pinchasi pinchasi later improved this to $\lfloor 5.5n\rfloor-3$ pinchasi . A construction in Fig. 12 (taken from horizon , just slightly enhanced) shows that the latter bound is almost optimal. Note the similarity with the construction in Fig. 10 (and also Fig. 18 further in the text).
@@
The final refinement in the upper bound comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq. must be non-maximal. Such analysis also gives an algorithm to actually construct a seq. of length $\mathscr{U}(n)$ which does not violate combinatorics of a regular union of $n$ $\triangle$ whose sides are segments of pseudolines.
For the starters, let us return to the example $\mathcal{L}=123214546417871$. It is maximal ($\mathbf{M}$ for short). Now, shift our attention to the cyan triangle. Or, better, the overlined fragment in (1). The produced seq. is also M. We argue that this extension to $\mathcal{L}$ can be reproduced from $\mathcal{L}$ itself. Indeed, starting from $\mathcal{L}$, remove all (but the last) $1$’s, and place $2$ in the end. This results in $232454647871\ldots 2$. For this to be $\mathbf{M}$, dots should be replaced by two numbers: $\mathcal{L}_{2}=232454647871ab2$. We assert that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$. The reasoning will now certainly look familiar. Let us methodically remove the last (reading left-to-right) symbol to appear, and merge its two neighbours. Such merge must always occur:
@@
This gives the only possible reconstruction: $\mathscr{S}(\mathcal{L})=\mathcal{L}_{2}=232454647871742$. The overlined fragment in (1) confirms it. Let us iterate this procedure (reconstructed symbols are printed in red):
@@
This exactly (!) corresponds to actual seq. extracted from (1). So far, all of them are $\mathbf{M}$. We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}$ by increasing all terms by $1$. This turns out to be a general rule.
Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences drawn from (1). For a while, $\mathbf{M}$ seq. continue. However, at a step $17$ (doubly underlined), a non-M sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs. Then it is followed with a $\mathbf{M}$ $2345436768631{\color[rgb]{1,0,0}3}2$. Next $2$ seq. are also $\mathbf{M}$. Then comes another shorther $56768612321645$ (length $14$), after that comes $\mathbf{M}$ $6768612321645{\color[rgb]{1,0,0}4}6$. Finally, after two more steps, everything closes up at the initial $1232145464178{\color[rgb]{1,0,0}7}1$. The conclusion: process almost always runs smoothly (and so the next seq. is deterministic), only during $24$ steps two glitches occur. The total number of edges in the $\bigcup$ equals the number of red symbols added, plus $6\cdot 8$ (the amount of black symbols).
If the reader has already decided that this is too confusing to follow, this is more than true. Seeing such an arbitrary-looking collection of numbers in (3), it is impossible to fathom what is actually happening. In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of a regular $(n+1)$-gons (first discovered by D. Roselle roselle ; see also sharir , p.250-232), we will see that… almost nothing is happening!
To witness, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$. First, divide it into $9$ decreasing portions as follows: | 1| 2| 3| 43| 53| 6| 76| 86321| 91||\,1|\,2\,|\,3\,|\,43\,|\,53\,|\,6\,|\,76\,|\,86321\,|\,91\,|. Given regular $(n+1)$-gon. For each vertex $i+1$, connect it to all vertices in a portion which starts from $i$. This way Fig. 13 is obtained. It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon with a marked vertex and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after renaming or reordering initial set of symbols).
@@
Figure 14: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 11; 1-9 also correspond exactly to seq. in (3).
Figure 15: Initial triangulations $\mathcal{T}_{0}$ of shifts with maximal weight $\omega(\mathcal{T}_{0})=12(n-1)-\gamma(n+1)$.
In our case, this triangulation convention is used for the natural order $1,2,\ldots,n$. If the order is $i,i+1,\ldots,i-1$, we adopt the following rule. First, rotate symbols in the seq. cyclically, so that the natural order is restored. This way, for $\mathcal{L}_{2}=232454647871742$ one gets $121343536768631$. Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to the position $i$.
Let as apply this rule to grasp what actual seq. extracted from (1) embody. The result is presented as Fig. 14. It is simultaneously surprising and relieving: nothing changes going from seq. 1 ($\mathcal{L}$) to seq. 2 ($\mathcal{L}_{2}$), except that the distinguished vertex moved one position. In light of such triangulation correspondence, the fact that the map $\mathscr{S}$ is well-defined becomes a tautology. Given a full triangulation of a regular $(n+1)$-gon. Then, if all edges ending at one particular vertex are removed, there is only one way to complete the triangulation by adding edges starting from this very vertex. Sounds like a joke, though it is not.
Moving on, such pattern (that is, the distinguished vertex jumps one ahead at a step) continues up to seq. 8. Going to seq. 9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction, and the distinguished vertex jumps two vertices ahead. This becomes clear if given some thought.
Now, what happens when a glitch (a shorter seq.) occurs? How it is succeeded by a $\mathbf{M}$ seq.? The same Fig. 14 (in the traditions of Intelligencer) might be considered as a proof without words of the following fact: at step $16\rightarrow 17$ (distinguished is vertex 8), one edge ending with $8$ is removed (at the same time, since the vertex changes from $8$ to $1$, a one-step rotation in positive direction occurs). At step $17\rightarrow 18$ (vertex 1) and edge with endpoint $1$ is added. After two more steps, another such incident takes place: one edge is removed, a new edge is reinstalled in the next iteration. As a final acumen to understand the whole procedure, note the following. When a vertex $8$ is replaced by $1$, instead of a rotation and a jump by $2$ vertices ahead, we can imagine that no rotation occurs, and the jump is an ordinary one, just by $1$ vertex. This way, after $24$ steps we arrive at vertex $7$. From its perspective, we witness the same initial triangulation.
Figure 17: Geometrically perfect polygon $\bigcup$ of inequivalent combinatorics. Top: 55 hexagons (type 1). Bottom: 77 quadrileterals (type 2). The maximal ratio of edge-lengths is shown. This number is close to the optimal value.
We come to understand that the whole process runs as follows. Start from a distinguished vertex $1$ in a triangulation of a regular $(n+1)$-gon. Perform $3n$ moves. An $i$th move amounts to removing few diagonals at this vertex; adding some other diagonals at the same vertex; counting their total number $t_{i}$; passing to the neighbour ahead. The last vertex where the move is performed is n2n-2 (afterwards, the jump to n1n-1 is made, yet followed by a halt). The goal is to obtain the same triangulation, only rotated by 3-3 vertices. The total combinatorial length of such $\triangle$ $\bigcup$ is then $\sum_{i=1}^{3n}t_{i}+6n$.
But this is exactly the definition of a triangulation shift! If no deletions are necessary, the triangulation coincides with itself rotated by 3-3 moves. The only two non-trivial examples exist for $n+1=6$ and $n+1=9$ (see Fig. 15). These exactly correspond to the Pentastar and the Octastar. For a generic $n+1$, if no deletions were necessary and we were to make the full $3$ turns, the weight would be 12n1212n-12. Define an auxiliary function $\gamma(\mathcal{T}_{0})=12n-12-\omega(\mathcal{T}_{0})$. It measures, so to say, the defect of a triangulation. It approximately equals the number of diagonals not shared by a 3-3 rotation of $\mathcal{T}_{0}$, plus a number of diagonals incident to some three consecutive vertices. That is, $\gamma(n)\sim n/3$. To precisely determine $\gamma$, we employ Python to do brute force calculations for $5\leq n\leq 16$. Fig. 15 shows initial values and champions (starting from $n+1=7$, not unique) for the corresponding values of $\gamma$. After the pattern is found, it is a standard routine to demonstrate that for $n\geq 5$, $\gamma(n+1)=n+2-2\lfloor\frac{n+1}{3}\rfloor$.
@@
Take $7$ linear univariate functions defined on the real line. Label them $1$ through $7$. Let $f$ be defined as the minimum of these. If, while taking the minimal value, we jump from one line to the other, it is obviously no longer possible to go back to the abandoned one. Therefore, the support sequence (telling which function provides the minimum as one progresses along the $x$-axis) for $f$ consist of at most $7$ symbols.
@@
However, if each linear function is defined on its own interval so that their $\bigcup$ is also an interval, the support sequence can be longer. This is clear. What is astonishing that it can be much, much, longer. As an example, consider $7$ segments as in Fig. 16. The support seq. $1213145416164653737572767$ has length $25$. What properties does such seq. have? Similarly as with $\mathcal{L}$, no two adjacent symbols are equal. At the same time, we immediately notice its subseq. like $1313$, $1414$, $3535$, $4646$, etc. This is a major distinction. We leave it as an exercise for the reader to prove that though subseq. abababab may be present, subseq. of the for ababaababa are not to be found. This is exactly the definition of a $(3,n)$-Davenport-Schintzel seq.
@@
The foundational part of the monograph by Sharir and Agarwal (sharir , p. 12-42) is devoted to investigating $(3,n)$-DS seq. The story how their true complexity unravelled is inspiring. The final answer – the maximal length of such seq. is $2n\alpha(n)+O(n)$, a result by Seth Pettie pettie – tells little about the story. In the preface to the book it is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
@@
Here is another major result. Consider a $(3,7)$-DS, now of maximal possible length $27$ (see seq. A002004 in OEIS): $121314151616543272737475767$. It can be realized as a lower envelope of $7$ pseudosegments (connected parts of $7$ pseudolines). Can it be realized as a lower envelope of $7$ segments? In other words, can pseudosegment construction be stretched? I do not know. What is certain that there is little hope (and most likely, untrue; see sharir , p. 112) that a certain maximal $(3,n)$-DS seq. for any sufficiently large $n$ can serve as a lower envelope of $n$ segments. However, Wiernik and Sharir proved that there exist $(3,n)$-DS seq. whose length asymptotically differs by a factor of at most $2$ from the length of the longest such seq. and which does serve as lower envelope of $n$ segments. Speaking philosophically, this shows that the Ackermann function, first discovered in logic and computability analysis, is intricately present in nature, if we assume that straight segments are. Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is what carries mathematics well outside the domain of natural sciences.
@@
To finish the construction, consider Fig. 16 once again. If we imagine each segment completed to a triangle by a left leg leaning slightly to the right and the right leg leaning slightly to the left, we clearly can get a simple polygon. Thus, finding $i(n)$ rests on other fundamental problems about $\bigcup$ and stretchability.
Figure 18: A construction showing that $c(n)\geq 11n-15$ ($n=7$ in the example; improvement to 11n1311n-13 is immediate). Without the orange $\triangle$, the union becomes regular, and this construction yields $r(n)\geq 10n-10$ (also, 10n810n-8 can be achieved).
Figure 19: A handwritten note made by Andrzej Schinzel in Vilnius, 2011.
Open (but closable) problems
Seq. $c(n)$ (vertices of all $\triangle$ are on the unit circle) and $r(n)$ (additionally, $\bigcup$ is regular) were not dealt with in this paper. Since $r(n)\leq e(n)$, this seq. is, obviously, bounded by a linear function. For $1\leq n\leq 8$ the equality holds. This might be the end of a friendship: numerical search produces cyclically arranged 99 $\triangle$ on the unit circle with $90$ edges, but no example with $91$ edges was found. On the other hand, Fig. 18 (without an orange $\triangle$) gives a construction showing that $r(n)\geq 10n-10$. Two black curves are circle arcs. By pushing a red $\triangle$ towards a brown one and a violet one towards a blue, a reflex angle at vertex $v$ can be cut twice. This raises the lower bound to 10n810n-8. There is, however, some space for further improvement.
Out of the four, the sequence $c(n)$ clearly stands out. Sure, $r(n)\leq c(n)\leq i(n)$. Fig. 18 gives a construction showing that $c(n)\geq 11n-13$. Yet, at the moment it seems that nothing substantially more can be said. The lower envelope technique does not work here. Moreover, Agarwal and Sharir plainly pose a problem about the maximal complexity of the lower envelope of line segments whose both ends are anchored at a circle arc (sharir , p. 112).
Finally, triangulation shifts is a tool not specific to $\triangle$ $\bigcup$. It also helps to determine $e_{p}(n)$, a maximal number of edges a regular $\bigcup$ of $n$ $p$-gons can have. In the definition, just replace $3n$ moves with $pn$ moves, and a rotation by 3-3 with a rotation by p-p. Sure, such general problem feels now too messy and premature. Yet, here is its sub-problem of independent interest.
A perfect proper $\bigcup$ of $n$ $p$-gons must satisfy $(n+1)e_{p}(n)=(4n-2)pn$. Each of them must correspond to a triangulation of a regular $(n+1)$-gon which, if rotated by p-p vertices, coincides with itself. This gives three basic types (in case $p=3$ represented by, respectively, the Star of David, the Pentastar, the Octastar). Type $1$: n=p1n=p-1, the number of edges M=2(p1)(2p3)M=2(p-1)(2p-3). Type $2$: n=2p1n=2p-1, M=(2p1)(4p3)M=(2p-1)(4p-3). Type $3$: n=3p1n=3p-1, M=2(2p1)(3p1)M=2(2p-1)(3p-1). If such perfect union can be realized with straight lines and, moreover, has a $n$-fold rotational symmetry, we call it geometrically perfect. For $p=6$, $n=5$ (type 1), and for $p=4$, $n=7$ (type 2) there exist two combinatorially different perfect unions. This corresponds to that fact that there exist two inequivalent (with respect to an action of a dihedral group $\mathbb{D}_{6}$) triangulations of a regular hexagon, and also two inequivalent (with respect to $\mathbb{D}_{8}$) triangulations of a regular octagon with a point-symmetry. All four can be realized geometrically (Fig. 17). However, in the last example, the shortest edge should be really short. It is plausible that there exists only a finite number of geometrically perfect proper polygon $\bigcup$. If so, this might be a queer addition to other finite catalogues of geometric objects ($5$ Platonic solids, $11$ Archimedean tilings, $15$ types of pentagonal tilings, and so on).
To round up, the author is happy to pass a torch to any interested reader. The first four of these questions might be immediately jumped upon.
  • 1)
@@
  • 2)
    Prove rigorously that $r(9)=90$.
  • 3)
    Improve the lower bound for $r(n)$.
  • 4)
    Find a perfect $\bigcup$ which is not realisable as geometrically perfect.
  • 5)
    What arguments are in favour and against the possible linear growth of $c(n)$?
  • 6)
    Can e(n+1)e(n)=14e(n+1)-e(n)=14 occur?
Since $\mathscr{U}(n+1)-\mathscr{U}(n)$ from $n=5$ onwards periodically repeats the pattern $11,11,13$, the affirmative answer to the last question is not ruled out.
@@
During the Journées Arithmétiques 2011 conference held Vilnius, Lithuania, Artūras Dubickas discussed a certain problem with Andrzej Schinzel. As was his custom, Prof. Schinzel later returned with a handwritten note containing the solution (see Fig. 19 containing the statement; afterwards comes the solution). Prof. Dubickas kindly presented this note to me. Though it is a bit of a stretch, but numbers $n=1,2,5,8$, giving rise to perfect triangle unions, are also in the Fibonacci sequence.